Cupcake123456:

Katrina is asked to simplify the expression -3a-4b-2(5a-7b) @ultrilliam @vocaloid @zarkam21

Vocaloid:

start by distributing this part 2(5a-7b) = ?

Cupcake123456:

(5a - 7b)2?

Vocaloid:

A(B-C) = A*B - A*C following the same logic 2(5a-7b) = ?

Vocaloid:

let's try and dissect A(B-C) = A*B - A*C to see what it means we take the number on the outside, A, and multiply it by the first number on the inside to get A*B then we take the number on the outside, A, and multiply it by the second number on the inside, to get A*C then the solution becomes A*B - A*C

Vocaloid:

applying this logic to 2(5a-7b) = ? we get 2*5a - 2*7b = ?

Cupcake123456:

where not supposed 2 b solving this were supposed to figure out which property

Vocaloid:

oh in that case she would start by ~distributing~ 2(5a-7b) so which property is this?

Cupcake123456:

-3a-4b-2(5a-7b)=-3a-4b-10a+14b

Vocaloid:

yes, so she ~distributed~ -2(5a-7b), what property is this?

Cupcake123456:

witch property is that ?

Vocaloid:

~distributed~

Vocaloid:

= distributive property

Cupcake123456:

the associative property

Cupcake123456:

?

Vocaloid:

emphasis on the word ""distributed"" = "distributive property"